Consider an annual coupon bond with a face value of ​$100​, 15 years to​ maturity, and a price of $94. The coupon rate on the bond is 2​%.

If you can reinvest coupons at a rate of 5​% per​ annum, then how much money do you have if you hold the bond to​ maturity?

The total proceeds from holding the bond to maturity are

​$nothing.

​(Round to the nearest​ cent.)
Answer #1

Total proceedings = Face value + Earnings

We have calculate earnings by using FV function in excel formula.

Rate = 5%

Nper = 15

Pmt = 100*2% = $2

PV = 0

Earnings = FV(rate,nper,pmt,PV,type)

=FV(5%,15,2,0,0)

=$43.16

Therefore earnings = $ 43.16

Total Proceedings = $100 + $43.16 = $146.16